a hockey puck is whqcked across the surface if a level frozen lame tjat is 400 m wide. if friction is not significant ehat is the oycks soeed when it reaches the other aide if its initial soeed was 50 m/s
Answer #1

It is given that the friction is negligible which employs acceleration is zero.

v^2 = u^2 + 2*a*s

v^2 = 50^2 + 2*0*400

V = 50 m/s

Hence, initial speed = final speed = 50 m/s